### v4

# BiMix: Bivariate Data Mixing Law for Language Model Pretraining

Abstract
Large language models have demonstrated remarkable capabilities across various tasks, primarily attributed to the utilization of diversely sourced data. However, the impact of pretraining data composition on model performance remains poorly understood. This paper introduces BiMix, a novel bivariate data mixing law that models the joint scaling behavior of domain proportions and data volume in language model pretraining. BiMix provides a systematic framework for understanding and optimizing data mixtures across diverse domains. Through extensive experiments on two large-scale datasets, we demonstrate BiMix’s high accuracy in loss extrapolation (mean relative error <0.2%) and its generalization to unseen mixtures (R2 >0.97). Optimization of domain proportions yields superior model performance compared to existing methods. Furthermore, we establish entropy-based measures as efficient proxies for data mixing, offering a computationally lightweight strategy. Our work contributes both insights into data mixing dynamics and practical tools for enhancing training efficiency, paving the way for more effective scaling strategies in language model development.
1 Introduction
Large language models (LLMs) have achieved remarkable success, revolutionizing capabilities for comprehending and generating human-like text across diverse applications, from question answering to code generation (Bubeck et al., 2023; OpenAI, 2024). As these models scale up, the composition of pretraining data becomes increasingly crucial for performance and generalization (Longpre et al., 2023). The emergence of multi-source datasets has presented both opportunities and challenges in LLM development (Gao et al., 2020; Shen et al., 2023; Chen et al., 2024), necessitating a deeper understanding of data mixing strategies.
Current approaches to data mixing often rely on heuristics (Touvron et al., 2023; Shen et al., 2023) or computationally expensive optimization techniques (Du et al., 2022; Xie et al., 2023; Fan et al., 2023). While these methods have shown promise, they require significant computational resources and lack a general framework for understanding the scaling behavior of mixed-domain training. The absence of a systematic approach to data mixing hinders efficient resource allocation and limits the ability to predict model performance across varied data compositions. Recent efforts have explored related techniques (Xia et al., 2024; Albalak et al., 2023; Shen et al., 2023), yet comprehensive and efficient solutions remain elusive.
The fundamental challenge lies in the complex interplay between different data domains in multi-source datasets. Existing research (Kaplan et al., 2020; Hoffmann et al., 2022) primarily focuses on scaling laws for individual metrics, overlooking this crucial aspect. This oversight hampers the progress towards more versatile models capable of excelling across multiple domains (Dong et al., 2024). A systematic mixing law remains to be developed to efficiently assess the importance of diverse data sources and understand their impact on model performance.
To address these challenges and fill the gap in current research, we introduce BiMix, a novel bivariate data mixing law that provides a systematic framework for understanding and optimizing data mixtures in LLM pretraining. Our approach is rooted in the observation that the scaling behavior of LLMs can be disentangled into two key components: domain mixing proportions and training data quantity (embodied by model training steps). By mathematically formulating the relationship between these components and model performance, BiMix offers a powerful tool for predicting and optimizing training outcomes.
We validate the proposed mixing law on two large-scale datasets, demonstrating its applicability across various scaling scenarios. Our experiments show that BiMix not only provides accurate predictions of model performance across different data mixtures but also enables optimization of domain proportions, outperforming existing high-cost methods in terms of convergence speed and downstream task performance.
The key contributions of this work are summarized as follows:
  • We introduce a novel bivariate mixing law BiMix for language model pretraining, which reveals the predictable effects of both data quantity and mixing proportions, providing significant interpretability and extensibility.
  • We offer insights into efficient mixture optimization, demonstrating that entropy measures serve as effective proxies, and provide practical guidance for data mixing.
  • We validate the utility and precision of BiMix through comprehensive experiments, highlighting its effectiveness in predicting and optimizing model performance across diverse datasets and training scenarios.
2 Related Work
Pretraining Data Mixtures  The coverage and diversity of pretraining data play significant roles in shaping the generalization capabilities of language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023). Data mixtures from multiple sources, such as the Pile (Gao et al., 2020) and ROOTS (Laurençon et al., 2022), are typically curated based on manually devised rules. However, the heuristics lack universal standards and portability. The GLaM dataset (Du et al., 2022) determined domain weights based on the component performance in a small model; however, specific details are not disclosed. SlimPajama-DC (Shen et al., 2023) investigated the effects of data mixtures using a set of predefined configurations and delivered several insights. Recently, DoReMi (Xie et al., 2023) and DoGE (Fan et al., 2023) proposed learning-based methods to optimize domain proportions by iterating between training reference and proxy models. These methods provide viable pathways but require considerable computational costs. In contrast, our study demonstrates that entropy proxies can produce data mixtures of comparable or even superior quality, while providing a more practical training-free solution. Besides, Chen et al. (2023) explored the effects of data sequencing from a curriculum learning perspective, whereas our research focuses on the concurrent integration of diverse data domains.
Neural Scaling Laws  Investigations into the scaling behavior of neural models have spanned across domains such as computer vision (Klug & Heckel, 2023; Zhai et al., 2022; Jain et al., 2023; Sorscher et al., 2022) and natural language processing (Ivgi et al., 2022; Gordon et al., 2021; Ghorbani et al., 2022; Bansal et al., 2022).  Kaplan et al. (2020) thoroughly evaluated the scalability of Transformer architectures across a wide range of model sizes and data volumes. Chinchilla (Hoffmann et al., 2022) identified similar scaling patterns through rigorous experimentation and suggested a slightly different configuration for compute-optimal pretraining. The impactful GPT-4 model (OpenAI, 2024) validated the predictive accuracy of scaling laws and underscored their important role in the development of large language models. Concurrently, additional research efforts seek to elucidate the principles governing scaling laws (Sharma & Kaplan, 2022; Michaud et al., 2023) and to investigate scaling effects on downstream tasks (Tay et al., 2022; Isik et al., 2024; Caballero et al., 2023; Cherti et al., 2023). In the context of data mixtures, Ye et al. (2024) proposed a composite exponential law to capture the interactions among domains; yet, its scalability is challenged by increased complexity for expanding domain numbers, as compared in Appendix C. Our study is distinguished by two key aspects: First, we introduce a scalable mixing law that accurately captures the scaling behavior associated with the composition of training datasets, demonstrating the modeling ability to up to 22 domains. Second, the proposed bivariate mixing law jointly models two input variables, domain proportion and data volume, thereby offering broader applicability.
3 The Proposed BiMix
Existing scaling law research primarily investigates the variation of a single scalar metric related to trained models concerning certain scaling factors. A prominent example is the relationship between a model’s validation loss and the amount of parameters or training tokens. However, in practice, the training datasets for large language models encompass diverse data domains, and these scaling laws only capture the predictability of the averaged validation loss across multiple domains. Since each domain corresponds to vastly different corpora, knowledge, and formats, modeling solely the average provides a coarse estimate of the model’s performance and fails to reflect the predictability of individual domains.
In the context of data-centric language modeling, this study examines the scaling behavior of pretrained models across finer-grained data domains. Notably, simply applying existing scaling laws to each domain is inappropriate, as the amount of training data for one single domain is not an independent variable; rather, it is determined by the total amount of training data and the proportion allocated to that domain, calculated as $|\mathcal{D}_{i}|=|\mathcal{D}|\times r_{i}$. The training data for a specific domain can be adjusted either by changing the total training data or by modifying the allocated proportion. Consequently, the data mixing modeling inherently involves a bivariate joint effect. Moreover, when the total amount of training data is fixed, any change in the proportion allocated to one domain will also affect the proportions (and thus the training amounts) of all other domains. This interdependence among domains highlights the need for dedicated research on the scaling laws of data mixing.
3.1 Formulation
Assume the training dataset consists of $m$ domains, each allocated a proportion $r_{i}$ under the unit-sum constraint:
$\accentset{\varrightharpoonup}{r}=(r_{1},r_{2},\ldots,r_{m})\quad\text{subjectto }\sum_{i=1}^{m}r_{i}=1.$ (1)
We propose BiMix as a system of equations to model the vectorized scaling laws of data mixing across the $m$ related domains in terms of the domain proportions $\accentset{\varrightharpoonup}{r}$ and training steps $s$:
$\accentset{\varrightharpoonup}{L}(\accentset{\varrightharpoonup}{r},s)=\left(L_{1}(r_{1},s),L_{2}(r_{2},s),\ldots,L_{m}(r_{m},s)\right).$ (2)
In this context, the validation loss of the model on the $i$-th domain, given the domain’s proportion and the number of training steps, is defined by the following function:
$L_{i}(r_{i},s)=\frac{A_{i}}{r_{i}^{\alpha_{i}}}\left(\frac{B_{i}}{s^{\beta_{i}}}+C_{i}\right)\quad\text{for }i=1,2,\ldots,m,$ (3)
where the constants $A_{i}$, $B_{i}$, $C_{i}$ and the exponents $\alpha_{i},\beta_{i}$ are coefficients to be fitted. Notably, only five fitting coefficients per domain are needed to capture the joint scaling behavior concerning both the mixing proportion and the total training volume. This linear scalability with the number of domains offers a significant advantage over the quadratic complexity of other modeling approaches, substantially reducing the number of observational data points required for fitting. A detailed discussion of the complexity analysis can be found in Appendix C.
3.2 Observing Scaling Behaviors by Disentangling Variables
It is important to recognize that scaling laws are fundamentally empirical formulas, providing mathematical descriptions that closely approximate real-world scaling phenomena. The construction of our proposed Eq. 3 is informed by a disentangled observation of the scaling behaviors of two variables, designed to offer strong interpretability. Next, we will elaborate on the observed scaling behaviors along with intuitive visualizations from the perspectives of the two input variables, as well as the considerations that guided the derivation of the final functional form.
(a) The Pile
(b) SlimPajama
Figure 1: Visualization of the fitting results for Eq. 3 at different domain proportion values, showing the relationship between validation loss and training steps. Each subplot corresponds to a specific domain within different datasets; the points represent the actual observed validation loss, while the dotted lines indicate the fitted results. Both axes are on a logarithmic scale.
Scaling Training Steps Under Fixed Domain ProportionsFigure 1 showcases the scaling behavior of the input variable $s$ on the Pile and SlimPajama datasets. Both the x and y axes are visualized on a logarithmic scale. Each line in the subplots corresponds to a specific value of the current domain proportion $r_{i}$, and the six lines represent the outcomes of six model training sessions on different data mixtures. These lines illustrate how the validation loss of each domain changes as the training steps increase. The discrete points in the figure indicate the actual evaluation results of the trained model, collected every 5 billion training tokens, while the dotted lines represent the fitted Eq. 3 derived from these observational data points. Overall, it is clear that the proposed mixing law fits the observed data points closely, and the curves exhibit a consistent pattern of downward curvature. In research related to scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; OpenAI, 2024), the average validation loss of the model is typically described as following a power law with an irreducible term in relation to the training steps. The decline pattern observed across the various domains aligns with this behavior and the loss for $i$-th domain can be expressed as:
$L_{i}(s\mid r_{i})=\frac{\tilde{B}_{i}}{s^{\tilde{\beta}_{i}}}+\tilde{C}_{i}.$ (4)
Here, $\tilde{B}_{i}$ and $\tilde{\beta}_{i}$ are the scaling factor and exponent of the power-law function, while the additional deviation term $C_{i}$ is understood as the lower bound for language modeling (Bishop, 2006; Henighan et al., 2020). Further examination of the curves in each subplot reveals that they approximately exhibit a shifting relationship in logarithmic space, described by:
$\log L_{j}(s\mid r_{j})=\log L_{i}(s\mid r_{i})+\log F_{j}=\log(L_{i}(s\mid r_{i})\cdot F_{j}),$ (5)
where $\log F_{j}$ represents an offset constant. This means that the loss function $L_{j}$ of $j$-domain can be obtained by multiplying $L_{i}$ by a scaling factor $F_{j}$:
$L_{j}(s\mid r_{j})=L_{i}(s\mid r_{i})\cdot F_{j}$ (6)
Extending this relationship to all curves within the same domain, the constant $F_{j}$ can be transformed into a function $f$ related to $r_{i}$, which is applied to a base scaling function of training steps:
$L_{i}(s\mid r_{i})=L_{\text{base}}(s\mid r_{i})\cdot f(r_{i}).$ (7)
Combined with Eq. 4, this multiplicative decomposition is clearly associated with Eq. 3, with the following relations: $\tilde{B}_{i}=B_{i}\cdot f(r_{i})$, $\tilde{C}_{i}=C_{i}\cdot f(r_{i})$, and $\tilde{\beta}_{i}=\beta_{i}$.
(a) The Pile
(b) SlimPajama
Figure 2: Visualization of the fitting results for Eq. 3 at different numbers of training steps, showing the relationship between validation loss and domain proportion. Each subplot corresponds to a specific domain within different datasets; the points represent the actual observed validation loss, while the dotted lines indicate the fitted results. Both axes are on a logarithmic scale.
Scaling Domain Proportions Under Fixed Training Steps  From the other perspective, we analyze how changes in the proportion of a single domain affect its validation loss. The visualization in Fig. 2 follows a similar setting as before. Each line in the figure represents the relationship between validation loss and domain proportion at a specific number of training steps (i.e., training data volume). The points indicate actual observed values, while the dotted line represents the fitted results of Eq. 3. The most notable difference from the previous visualization is the prominent linear relationship observed. This straight line on a logarithmic scale strongly supports the standard power-law function. Thus, the pattern represented by a single straight line in the figure can be expressed as:
$L_{i}(r_{i}\mid s)=\frac{\tilde{A}_{i}}{r_{i}^{\tilde{\alpha}_{i}}}.$ (8)
The collection of straight lines within each subplot can be shifted relative to one another in logarithmic space, leading to the following derivation:
$\log L_{j}(r_{j}\mid s)=\log L_{i}(r_{i}\mid s)+\log G_{j}=\log(L_{i}(r_{i}\mid s)\cdot G_{j}).$ (9)
This implies:
$L_{j}(r_{j}\mid s)=L_{i}(r_{i}\mid s)\cdot G_{j}$ (10)
Considering the unified modeling of these straight lines, the constant $G_{j}$ can be converted to a function $g$ , yielding the following relationship:
$L_{i}(r_{i}\mid s)=L_{\text{base}}(r_{i}\mid s)\cdot g(s).$ (11)
Relating this with Eqs. 9 and 3, we obtain the following mappings $\tilde{A}_{i}=A_{i}\cdot g(s)$ and $\tilde{\alpha}_{i}=\alpha_{i}$.
Remark Through the disentanglement of the two input variables, we have identified a separable scaling effect between domain proportions $r_{i}$ and the number of training steps $s$. Given that dual multiplicative weighting functional forms were derived from both perspectives, we integrated them to construct a mutually modulated bivariate mixing law that encompasses strong interpretability.
4 Experimental Setup
Datasets  We employed two well-recognized datasets spanning diverse domains to conduct comprehensive experiments. The Pile (Gao et al., 2020) is a diverse language modeling dataset comprising 22 subsets with a total of 825 GiB of textual data. SlimPajama (Shen et al., 2023) is a high-quality, seven-domain dataset that has been rigorously deduplicated and refined to 627B tokens from the extensive 1.2T RedPajama dataset (Together Computer, 2023). Following the preprocessing procedures in Xie et al. (2023), we packed all samples within each domain and chunked them into sequences of 1024 tokens for improved training efficiency.
Model Architecture  We employed decoder-only transformers based on the DoReMi architecture (Xie et al., 2023). The base model comprises 12 decoder blocks, each with 768-dimensional embeddings, 12 attention heads, and 4×4$\times$4 × MLP hidden size, matching the DoReMi 280M specifications. For scaled-up experiments on optimized mixtures, we expanded to 16 blocks with 2048-dimensional embeddings and 32 attention heads, aligning with the DoReMi 1B model. All models use the GPT-NeoX tokenizer (Black et al., 2022) with a vocabulary size of 50,277.
Training Details  Experiments were conducted under controlled hyperparameters. Each training run consisted of up to 200,000 update steps with a global batch size of 512. The optimizer used was AdamW (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1\times 10^{-8}$, and a weight decay of 0.01. The learning rate, initialized at $1\times 10^{-3}$, decayed exponentially by a factor of 10×10$\times$10 × over the course of training. We leveraged data parallelism across eight NVIDIA A100 80GB GPUs and bfloat16 mixed precision to improve training throughput. As a reference for the training cost, a single round of DoReMi training took 670 GPU hours on this infrastructure.
Fitting Details  As noted previously, we trained models for up to 200,000 update steps on each data mixture, which corresponds to approximately 100 billion tokens. Model evaluation results were collected every 5 billion tokens, allowing for a maximum of 20 assessments, denoted by symbol $n$. During the training of the model on a given data mixture, validation losses for all domains can be obtained simultaneously, with $m$ representing the number of domains. When experiments are conducted on $k$ different data mixtures, we can potentially gather up to $m\times n\times k$ data points in the form of tuples $\langle r_{i},s,L_{i}\rangle$ for fitting BiMix. For the Pile dataset, $m=22$, while for SlimPajama, $m=7$. Upon collecting the observational data points, we employ the Trust Region Reflective algorithm (Branch et al., 1999; Virtanen et al., 2020) to fit the coefficients in Eqs. 2 and 3.
Evaluation Metrics  The investigated mixing law is fundamentally a specific form of scaling law, widely recognized for its ability to describe the predictability of model loss. Following the prevalent consensus in relevant literature, we primarily report the model’s validation loss on each domain, which is sometimes referred to as log-perplexity. To provide a more intuitive understanding of the model’s actual performance, we also evaluate model performance on downstream NLP tasks when necessary. The benchmarks from DoReMi are utilized for generative question-answering, specifically WebQuestions (Berant et al., 2013), LAMBADA (Paperno et al., 2016), and TriviaQA (Joshi et al., 2017). We employ the same one-shot prompting and Exact Match metric as used in DoReMi. A response is considered correct if and only if the characters of the model’s prediction exactly match those of the True Answer.
Candidate Mixtures  Estimating the coefficients in Eqs. 2 and 3 requires a series of observational data points. These are obtained by training on various candidate data mixtures for a limited number of iterations. We consider the following three types of mixtures:
  1. (a)
    Baseline: Represents the original proportions of the datasets, reflecting the intentions of the dataset creators or the inherent distribution of the data collection process.
  2. (b)
    DoReMi (Xie et al., 2023): This approach tunes domain weights by iteratively training reference and proxy models through group distributionally robust optimization. The tuned proportions for the Pile dataset were directly taken from the released results, while for the SlimPajama dataset, we strictly executed the official code to optimize the proportions.
  3. (c)
    Entropy: Measures that serve as efficient proxies for lightweight data mixing, including Shannon entropy (SE), conditional entropy (CE), joint entropy (JE), and von Neumann entropy (VNE). Specifically, we calculate the entropy metric for all samples in each domain to represent that domain’s data diversity or importance. These entropy values are then normalized across domains to yield mixing proportions that sum to 1. For example, in the case of conditional entropy, we first tokenize the original dataset, resulting in a token set $\mathcal{D}=(\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{m})$, where each domain $\mathcal{D}_{i}$ is a set of token sequences $\{(x_{1},x_{2},\dots,x_{T})\}$ of equal length $T$. The conditional entropy for one domain is computed as follows:
    $H_{i}(X_{i}^{(t+1)}\mid X_{i}^{(t)})=-\sum_{x\in X_{i}^{(t)}}\sum_{x^{\prime}\in X_{i}^{(t+1)}}P(x,x^{\prime})\log P(x^{\prime}\mid x),$ (12)
    where $X_{i}^{(t)}$ and $X_{i}^{(t+1)}$ are sets of tokens at positions $t$ and $t+1$, respectively. The joint probability $P(x,x^{\prime})$ and the conditional probability $P(x^{\prime}\mid x)$ are both statistical estimations on the token set. The mixing proportions $(r_{1},r_{2},\ldots,r_{m})$ are derived by exponentially normalizing the entropy measures:
    $r_{i}=\frac{e^{H_{i}}}{\sum_{j=1}^{m}e^{H_{j}}}.$ (13)
    The resulting proportions place greater emphasis on domains with higher entropy, indicating greater uncertainty, to enhance the learning process. Notably, implementing entropy measurement is highly efficient, as it can be seamlessly integrated into the tokenization process with negligible overhead. Details about the various entropy measures can be found in Appendix A.
5 Results and Analysis
We illustrate the applicability of BiMix from three dimensions: Section 5.1 validates its scalability regarding training data volume; Section 5.2 demonstrates the fitted law’s generalization across different mixtures; and Section 5.3 presents a direct method for optimizing domain proportions for application in larger-scale models. Finally, in Section 5.4, we compare data mixtures driven by different entropy measures, offering a streamlined and efficient strategy for data mixing.
5.1 Extrapolating Losses on Scaled-Up Data
Scaling laws are primarily used to extrapolate model metrics when scaling up the data volume. We held out the validation losses at the final 200,000 steps as the prediction target, and used the remaining observational data points to fit Eqs. 2 and 3. After fitting the coefficients, we used the law to predict the target loss for each domain. Assuming the ground truth loss is $y$ and the BiMix-predicted loss is $y^{\prime}$, we calculate the relative prediction error as $|y-y^{\prime}|/{y}$.
Table 1: Relative prediction error between the real validation loss and the BiMix-predicted validation loss for the final training step. Fitting is performed on all domains for each mixture, reporting the mean, worst, and best errors across domains.
Dataset Mixture Relative Prediction Error $\downarrow$
Mean (%) Worst (%) Best (%)
The Pile (22 domains) Baseline 0.16 0.43 0.03
DoReMi 0.19 0.95 0.02
CE 0.17 0.67 0.05
SlimPajama (7 domains) Baseline 0.18 0.29 0.14
DoReMi 0.17 0.24 0.10
CE 0.18 0.31 0.12
The results are presented in Table 1. It is evident that BiMix extrapolates losses with remarkable accuracy. For both the Pile and SlimPajama datasets, the mean relative error remains below 0.2%. Even for the worst-performing domain, the prediction error is less than 1.0% (DoReMi mixture on the Pile). Moreover, when comparing the error distributions across domains in the two datasets, it is observed that both the worst and best errors on SlimPajama are closer to the mean error, whereas the variance is higher on the Pile. This can be attributed to the greater diversity of the Pile, which comprises up to 22 domains and presents a more significant challenge. In contrast, the meticulous data deduplication applied to SlimPajama helps to reduce cross-domain interference and fitting noise.
5.2 Estimating Data Mixtures without Training
When training language models on multi-source data, determining the appropriate mixing proportions for different domains remains a persistent challenge for practitioners. The naive approach is trial and error, where a few random data mixtures are generated to train models and select the best-performing one. However, as both data and model scales continue to grow, each training session represents a significant expenditure; thus, training large language models has become a careful and strategic process. The proposed BiMix offers a cost-bounded solution based on scaling laws. By training models on a limited number of data mixtures and collecting sufficient observational data points to fit BiMix, we can estimate the effectiveness of any given data mixture without actually conducting training. This approach allows for prospective evaluation of candidate data mixtures before incurring substantial computational costs, helping to eliminate poor options and prioritize effective training configurations.
(a) The Pile
(b) SlimPajama
Figure 3: Correlation between the observed validation losses (x-axis) and the BiMix-predicted losses (y-axis) across training iterations with the Baseline and DoReMi mixtures.
We trained models on the four entropy-driven data mixtures and collected observational data points to fit BiMix. This fitted model was then used to predict the validation losses for each domain at each training step on the Baseline and DoReMi mixtures. As shown in Fig. 3, the data points from the top right to the bottom left depict the convergence direction as training iterations progress. The x-axis represents the actual loss observed during training, while the y-axis represents the loss predicted by the fitted BiMix. The presence of compact linear trends indicates a strong positive correlation between the two losses. In the lower-left corner of each subplot, the two stars represent the final losses at the end of model training. In Fig. 3(a), observing the y-axis reveals that the law predicts the final loss for DoReMi to be lower than that for Baseline, which is consistent with the actual relative magnitudes displayed on the x-axis; Fig. 3(b) demonstrates a similar effect.
Table 2: Goodness of fit measured by the coefficient of determination (R2) on validation mixtures. The better the fit, the closer the value is to 1.
Dataset Mixture Goodness of Fit (R2) $\uparrow$
Mean (%) Worst (%) Best (%)
The Pile (22 domains) Baseline 0.9748 0.7911 0.9974
DoReMi 0.9744 0.7864 0.9972
SlimPajama (7 domains) Baseline 0.9940 0.9896 0.9962
DoReMi 0.9945 0.9904 0.9970
Table 2 presents a quantitative assessment of the goodness of fit. We employ the commonly used coefficient of determination (R2(Wright, 1921) metric, which is computed based on the residual sum of squares $\text{SS}_{\text{res}}=\sum_{j=1}^{n}(y_{j}-y^{\prime}_{j})^{2}$ and the total sum of squares $\text{SS}_{\text{tot}}=\sum_{j=1}^{n}(y_{j}-\bar{y}_{j})^{2}$:
$\text{R}^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{tot}}},$ (14)
where $n$ is the number of different training volumes. The R2 value is computed over the sequence of $n$ data points for each domain, on a logarithmic scale to reduce deviation. The better the fit, the closer the value is to 1. Overall, the high mean R2 values suggest that BiMix exhibits good generalization across new data mixtures. The fitted law demonstrates a better fit on the SlimPajama dataset compared to the Pile, as reflected by higher mean and worst R2 values, aligning with the observations made in Section 5.1.
5.3 Optimizing Domain Proportions for Improved Performance
Recall that BiMix in Eq. 2 describes a system of related equations in which the input variable $\accentset{\varrightharpoonup}{r}$ adheres to a unit-sum constraint. This vectorized formulation facilitates the direct optimization of the proportions across various domains. Consider a common objective of minimizing the model’s average validation loss across domains, defined as:
$\bar{L}(\accentset{\varrightharpoonup}{r},s)=\sum_{i=1}^{m}L_{i}(r_{i},s),$ (15)
where $L_{i}$ is the loss function specific to $i$-th domain with proportion $r_{i}$ and training steps $s$. Solving the following constrained minimization problem yields an optimized domain proportion vector:
$\accentset{\varrightharpoonup}{r}^{*}=\operatorname*{arg\,min}_{r_{1},r_{2},\ldots,r_{m}}\bar{L}(\accentset{\varrightharpoonup}{r},s)\quad\text{subject to}\quad\sum_{i=1}^{m}r_{i}=1.$ (16)
This presents a classic constrained optimization problem, which can be addressed using Lagrange multipliers and numerical methods (Virtanen et al., 2020). For fitting and optimization, we utilized all observational data points from the four entropy-driven data mixtures, as they serve as effective proxies in practical applications (discussed in Section 5.4).
(a) The Pile
(b) SlimPajama
Figure 4: Comparison of average downstream accuracy of 1B models trained on different data mixtures. Details regarding specific tasks and the Exact Match metric can be found in Section 4.
To validate the efficacy of this approach, we adopt a strategy similar to that in DoReMi (Xie et al., 2023), whereby the optimized mixtures derived from smaller models are utilized to train a larger model with billion-level parameters. Figure 4 illustrates how downstream performance varies as the number of training steps increases. Recent advancements in mixture optimization are also included for comparison, including RegMix (Liu et al., 2024) on the Pile and DoGE (Fan et al., 2023) on SlimPajama. The models trained on the BiMix-optimized mixtures demonstrated performance advantages throughout the training process. While we note that RegMix achieved performance comparable to that of BiMix, it is crucial to highlight that our approach not only optimizes mixtures but also provides a mathematical model for understanding mixing behavior. A detailed comparison of performance across each task is included in Appendix B.
5.4 Entropy Measures as Efficient Mixing Proxies
(a) The Pile
(b) SlimPajama
Figure 5: Comparison of log-perplexity evaluations for models trained on different data mixtures.
To collect the observational data points necessary for fitting the coefficients of BiMix, we trained models on a series of entropy-driven data mixtures. Entropy essentially quantifies the uncertainty of data distribution, thereby reflecting the fitting difficulty of domains. We hypothesized that it could serve as a valuable proxy. Figure 5 compares the average log-perplexity evaluated on validation sets for models trained on different data mixtures. It is evident that all models trained on these entropy-driven data mixtures exhibit lower log-perplexity compared to the baseline, indicating that the models have learned statistical patterns from the data more effectively. This finding suggests that entropy measures are indeed efficient mixing proxies, facilitating the streamlined initial construction of pretraining datasets. Among these entropy-driven proxies, conditional entropy (CE) consistently demonstrates lower log-perplexity, thereby being regarded as the preferred candidate through experiments.
6 Discussion and Future Work
Scaling laws model the empirical behavior of model outcomes with respect to certain variables, typically effective within a limited observational range. The applicability of our proposed mixing law under extreme conditions is not guaranteed. In our experiments, domain proportions ranged from 0.0007 to 0.7256, with maximum training data capped at approximately 100B tokens. However, existing research suggests that loss predictability may extend to larger scales (Kaplan et al., 2020; Hoffmann et al., 2022). Our study adheres to settings aligned with relevant work, ensuring consistent domain inclusion during both model training and evaluation (Xie et al., 2023). When evaluating the trained model on new domains, the model’s generalization capability and the correlation between new and training domains become primary considerations. This aspect, extending beyond the basic mixing laws studied in this work, warrants dedicated exploration. Our work contributes to the limited research on mixing laws, aiming to establish a foundation for more comprehensive studies.
Reflecting on broader implications, our findings on mixing laws can assist practitioners in optimizing computational resource allocation, promoting advancements in economical and environmentally friendly AI development. This vision is also relevant to the rapidly evolving field of multimodal large models (McKinzie et al., 2024), where processing images, videos, or audio may consume significant computational power. Consequently, exploring the mixing of multimodal training data presents vast opportunities for enhancing model efficiency and performance. Future work could focus on extending our mixing law framework to multimodal contexts, potentially leading to more efficient and effective training paradigms for next-generation AI models.
7 Conclusion
This paper introduces BiMix, a bivariate data mixing law for language model pretraining. BiMix accurately models the joint scaling behavior of domain proportions and training volume, enabling precise loss extrapolation and generalization to different mixtures. Our experiments demonstrate its effectiveness in optimizing domain proportions, outperforming existing methods. Additionally, we show that entropy-based measures serve as efficient proxies for lightweight data mixing. By offering a nuanced understanding of data mixing dynamics, this research contributes to the development of more efficient large-scale language models and opens avenues for further exploration in data-centric machine learning.
Appendix A Entropy Proxies
Given a text dataset such as Pile and SlimPajama, we concatenate all samples within each domain and tokenize them into fixed-length sequences of 1024 tokens. During tokenization, we concurrently record the occurrence frequencies of all unigrams and bigrams, in preparation for computing Shannon entropy, joint entropy, and conditional entropy. The procedure for von Neumann entropy is slightly different: we employ the FastText (Grave et al., 2018) embedding model to map each text sample into a 300-dimensional vector, which will be used to compute their pairwise similarities.
Given a tokenized dataset $\mathcal{D}=(\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{m})$, where each domain $\mathcal{D}_{i}$ is a set of token sequences $\{(x_{1},x_{2},\dots,x_{T})\}$ of equal length $T$, the following entropy proxies are computed.
Shannon Entropy (SE)
$H_{i}(\mathcal{D}_{i})=-\sum_{x\in X_{i}}P(x)\log P(x)$, where $X_{i}$ is the set of all available tokens in domain $\mathcal{D}_{i}$ and $P(x)$ denotes the probability of observing token $x$. This proxy quantifies the expected information content associated with token appearances in the dataset, indicative of the corpus diversity.
Joint Entropy (JE)
$H_{i}(X_{i}^{(t)},X_{i}^{(t+1)})=-\sum_{x\in{X}_{i}^{(t)}}\sum_{x^{\prime}\in{X}_{i}^{(t+1)}}P(x,x^{\prime})\log P(x,x^{\prime})$, where $X_{i}^{(t)}$ and $X_{i}^{(t+1)}$ represent the sets of tokens at positions $t$ and $t+1$ across all sequences in domain $\mathcal{D}_{i}$, respectively. The joint probability function $P(x,x^{\prime})$ was statistically estimated by observing a token $x$ at position $t$, followed by a token $x^{\prime}$ at position $t+1$. This metric measures the average uncertainty associated with consecutive token pairs and highlights the sequential dependencies in the dataset.
Conditional Entropy (CE)
$H_{i}(X_{i}^{(t+1)}\mid X_{i}^{(t)})=-\sum_{x\in X_{i}^{(t)}}\sum_{x^{\prime}\in X_{i}^{(t+1)}}P(x,x^{\prime})\log P(x^{\prime}\mid x)$, with $X_{i}^{(t)}$, $X_{i}^{(t+1)}$, and $P(x,x^{\prime})$ as previously defined. The term $P(x^{\prime}\mid x)$ denotes the conditional probability of observing a token $x^{\prime}$ at position $t+1$ given the presence of token $x$ at position $t$. This measures the anticipated level of surprise when predicting the next token in a sequence, providing a clearer understanding of the text’s predictability and its complex linguistic structure.
Von Neumann Entropy (VNE)
In physics, the von Neumann entropy extends the concept of Gibbs entropy from classical statistical mechanics to quantum statistical mechanics. For a quantum-mechanical system described by a density matrix $\rho$, the von Neumann entropy for domain $\mathcal{D}_{i}$ is defined as
$H_{i}(\rho_{i})=-\operatorname{Tr}(\rho_{i}\log\rho_{i}),$ (17)
where $\operatorname{Tr}$ denotes the trace operation and $\log$ the matrix logarithm. Recent research has highlighted its utility in quantifying the diversity of datasets from a system perspective (Friedman & Dieng, 2023). For simplicity, we will drop the subscript $i$ in subsequent discussions, but keep in mind that all discussions pertain to a single domain. In the context of data mixture, we define $\rho$ as $K/N\in\mathbb{R}^{N\times N}$, where $N=|\mathcal{D}_{i}|$. The matrix $K$ is determined by a positive semi-definite kernel $k:{\mathcal{D}}\times{\mathcal{D}}\to\mathbb{R}$, such that $K_{jk}=k(v_{j},v_{k})$ and $k(v_{j},v_{k})=1$ for $1\leq j,k\leq n$, where $v_{j}$ and $v_{k}$ are the embedding vectors processed by FastText. In practice, the von Neumann entropy is computed through the eigenvalues of the density matrix $\rho$:
$H_{i}(\rho)=-\sum_{i=1}^{N}\lambda_{i}\log\lambda_{i},$ (18)
where the eigenvalues $\lambda_{i}$ represent the probability distributions akin to quantum states.
Appendix B Downstream Task Performance
(a) WebQuestions
(b) LAMBADA
(c) TriviaQA
Figure 6: Detailed downstream task performance of the 1B model trained with various data mixtures on the Pile dataset.
(a) WebQuestions
(b) LAMBADA
(c) TriviaQA
Figure 7: Detailed downstream task performance of the 1B model trained with various data mixtures on the SlimPajama dataset.
Figures 6 and 7 present the downstream performance of the 1B model trained on various data mixtures for the Pile and SlimPajama datasets. Key observations include:
  • BiMix-trained models consistently outperform others across most tasks.
  • Performance on WebQuestions and LAMBADA shows some variability, but BiMix models maintain overall superiority.
  • In TriviaQA, all models demonstrate a stable upward trend with increasing iterations, with BiMix mixtures showing clear advantages.
These results underscore the effectiveness of BiMix-optimized data mixtures in enhancing model performance across diverse downstream tasks.
Appendix C Complexity Analysis
Table 3: Complexity comparison of mixing laws.
Mixing Law Number of Coefficients
1 domain 1 target $m$ domains 1 target $m$ domains $n$ targets
$L_{i}(r_{1\dots m})=c_{i}+k_{i}\exp(\sum_{j=1}^{m}t_{ij}r_{j})$(Ye et al., 2024) $m+2$ $m^{2}+2m$ $m^{2}n+2mn$
$L_{i}(r_{i},s)={\frac{A_{i}}{r_{i}^{\alpha_{i}}}\left(\frac{B_{i}}{s^{\beta_{i}}}+C_{i}\right)}$  (BiMix) $2$ $2m$ $5m$
Table 3 compares the fitting complexity of our proposed BiMix with the concurrent composite exponential law by Ye et al. (2024). Both equations model the validation loss for multi-domain language modeling. The exponential law, $L_{i}(r_{1\dots m})$, operates on all domain proportions $(r_{i},r_{2},\dots,r_{m})$ without considering training steps, while our mixing law incorporates both domain proportion $r$ and training steps $s$. We analyze complexity across three scenarios, progressing from simple to complex.
Base Case: Fitting an Individual Domain.
The exponential law requires $m+2$ fitting coefficients, aggregating proportions across $m$ domains with $m$ weighting coefficients $t_{ij}$, plus scaling coefficient $k_{i}$ and translation coefficient $c_{i}$. Our bivariate mixing law simplifies to Eq. 8 for fixed training steps, needing only two coefficients: a scaling factor $\tilde{A}_{i}$ and an exponent $\tilde{\alpha}_{i}$.
General Case: Fitting All Domains.
For $m$ domains, both mixing law’s coefficient requirement scales by $m$, while our mixing law maintains an order of magnitude fewer coefficients.
Extensive Case: Fitting Across Multiple Targets.
The exponential law, not accounting for variable $s$, requires $(m^{2}+2m)n$ coefficients for $n$ training step targets. Our bivariate mixing law, incorporating both domain proportions and training steps, shares fitting coefficients across various training steps. Specifically, only five coefficients are required to fit an individual domain across different training steps; this number scales linearly to $5m$ when generalized to all $m$ domains.
Overall, Our bivariate scaling law consistently requires significantly fewer coefficients compared to the composite exponential law in Ye et al. (2024). This reduction translates to fewer required observations for effective fitting, enabling our law to be fit with just a few (potentially as few as two) candidate mixtures, whereas the exponential law needs tens of mixtures. The computational efficiency of our method offers economic and environmental benefits through reduced resource utilization, while simultaneously achieving better data mixtures and enhanced model performance, as demonstrated in Section 5.3.
Appendix D Comparision with Recent Mixing Law
(a) The Pile
(b) SlimPajama
Figure 8: Performance comparison of the 1B model trained on the SlimPajama dataset with recent work by Ye et al. (2024).
We extended the comparison to include the optimal data mixture identified by Ye et al. (2024) for training the 1B model on the SlimPajama dataset. Notably, the exponential law proposed by Ye et al. faces scalability challenges when applied to the more diverse, 22-domain Pile dataset due to the previously discussed quadratic complexity.
Analysis of the results presented in  Fig. 8 reveals several key findings:
  • Both our BiMix-optimized mixture and Ye et al. (2024)-optimized mixture demonstrated accelerated model convergence compared to the default Baseline mixture.
  • Our BiMix-optimized data mixture achieved equivalent log-perplexity to the Baseline using only 50% of the training steps required by Ye et al. (80,000 vs. 160,000), indicating more effective data utilization.
  • While the 1B model trained on the Ye et al.-optimized mixture showed performance comparable to the Baseline on downstream tasks, the model trained on our BiMix-optimized mixture exhibited substantial advantages.
These results further underscore the effectiveness of our BiMix approach in optimizing data mixtures for large language model training, offering both improved convergence speed and enhanced downstream task performance.
Appendix E Mixture Recipes
Tables 4 and 5 provide detailed compositions of the candidate data mixtures employed for the Pile and SlimPajama datasets, respectively. These mixtures form the basis of our experiments and are integral to the evaluation of our BiMix approach.
Table 4: Data mixtures on the Pile dataset
Domain Baseline DoReMi SE CE JE VNE
ArXiv 0.088 617 813 216 488 410.088617813216488410.088\,617\,813\,216\,488\,410.088 617 813 216 488 41 0.053 539 443 761 110 3060.0535394437611103060.053\,539\,443\,761\,110\,3060.053 539 443 761 110 306 0.033 564 845 50.03356484550.033\,564\,845\,50.033 564 845 5 0.034 977 458 70.03497745870.034\,977\,458\,70.034 977 458 7 0.023 615 029 251 884 10.02361502925188410.023\,615\,029\,251\,884\,10.023 615 029 251 884 1 0.025 779 460 90.02577946090.025\,779\,460\,90.025 779 460 9
BookCorpus2 0.004 369 643 887 833 8660.0043696438878338660.004\,369\,643\,887\,833\,8660.004 369 643 887 833 866 0.003 705 330 193 042 7550.0037053301930427550.003\,705\,330\,193\,042\,7550.003 705 330 193 042 755 0.034 897 088 50.03489708850.034\,897\,088\,50.034 897 088 5 0.046 073 6230.0460736230.046\,073\,6230.046 073 623 0.032 341 390 467 724 7040.0323413904677247040.032\,341\,390\,467\,724\,7040.032 341 390 467 724 704 0.021 039 950 60.02103995060.021\,039\,950\,60.021 039 950 6
Books3 0.072 011 395 173 209 360.072011395173209360.072\,011\,395\,173\,209\,360.072 011 395 173 209 36 0.075 663 350 522 518 160.075663350522518160.075\,663\,350\,522\,518\,160.075 663 350 522 518 16 0.054 401 6590.0544016590.054\,401\,6590.054 401 659 0.069 749 976 10.06974997610.069\,749\,976\,10.069 749 976 1 0.076 325 855 494 299 720.076325855494299720.076\,325\,855\,494\,299\,720.076 325 855 494 299 72 0.028 257 869 70.02825786970.028\,257\,869\,70.028 257 869 7
DM Mathematics 0.020 432 462 197 971 4470.0204324621979714470.020\,432\,462\,197\,971\,4470.020 432 462 197 971 447 0.001 869 944 739 155 471 30.00186994473915547130.001\,869\,944\,739\,155\,471\,30.001 869 944 739 155 471 3 0.004 740 5770.0047405770.004\,740\,5770.004 740 577 0.007 773 0940.0077730940.007\,773\,0940.007 773 094 0.000 741 208 763 868 702 90.00074120876386870290.000\,741\,208\,763\,868\,702\,90.000 741 208 763 868 702 9 0.043 383 888 40.04338388840.043\,383\,888\,40.043 383 888 4
Enron Emails 0.002 939 416 393 568 1880.0029394163935681880.002\,939\,416\,393\,568\,1880.002 939 416 393 568 188 0.003 953 147 213 906 050.003953147213906050.003\,953\,147\,213\,906\,050.003 953 147 213 906 05 0.041 631 948 30.04163194830.041\,631\,948\,30.041 631 948 3 0.027 176 178 60.02717617860.027\,176\,178\,60.027 176 178 6 0.022 757 758 155 987 6180.0227577581559876180.022\,757\,758\,155\,987\,6180.022 757 758 155 987 618 0.071 720 060 60.07172006060.071\,720\,060\,60.071 720 060 6
EuroParl 0.007 475 643 162 169 9140.0074756431621699140.007\,475\,643\,162\,169\,9140.007 475 643 162 169 914 0.011 973 137 035 965 920.011973137035965920.011\,973\,137\,035\,965\,920.011 973 137 035 965 92 0.068 138 717 90.06813871790.068\,138\,717\,90.068 138 717 9 0.031 448 426 10.03144842610.031\,448\,426\,10.031 448 426 1 0.043 103 027 821 376 750.043103027821376750.043\,103\,027\,821\,376\,750.043 103 027 821 376 75 0.070 945 6710.0709456710.070\,945\,6710.070 945 671
FreeLaw 0.040 288 640 046 477 530.040288640046477530.040\,288\,640\,046\,477\,530.040 288 640 046 477 53 0.038 014 464 080 333 710.038014464080333710.038\,014\,464\,080\,333\,710.038 014 464 080 333 71 0.037 513 087 50.03751308750.037\,513\,087\,50.037 513 087 5 0.039 405 403 30.03940540330.039\,405\,403\,30.039 405 403 3 0.029 734 037 757 738 0680.0297340377577380680.029\,734\,037\,757\,738\,0680.029 734 037 757 738 068 0.033 367 036 50.03336703650.033\,367\,036\,50.033 367 036 5
GitHub 0.055 383 585 535 041 760.055383585535041760.055\,383\,585\,535\,041\,760.055 383 585 535 041 76 0.032 475 057 989 358 90.03247505798935890.032\,475\,057\,989\,358\,90.032 475 057 989 358 9 0.038 158 005 80.03815800580.038\,158\,005\,80.038 158 005 8 0.045 077 537 50.04507753750.045\,077\,537\,50.045 077 537 5 0.034 598 815 296 213 770.034598815296213770.034\,598\,815\,296\,213\,770.034 598 815 296 213 77 0.102 148 413 40.10214841340.102\,148\,413\,40.102 148 413 4
Gutenberg (PG-19) 0.021 821 973 865 940 4350.0218219738659404350.021\,821\,973\,865\,940\,4350.021 821 973 865 940 435 0.029 194 029 048 085 2130.0291940290480852130.029\,194\,029\,048\,085\,2130.029 194 029 048 085 213 0.034 849 1780.0348491780.034\,849\,1780.034 849 178 0.053 462 684 10.05346268410.053\,462\,684\,10.053 462 684 1 0.037 476 512 874 596 460.037476512874596460.037\,476\,512\,874\,596\,460.037 476 512 874 596 46 0.021 066 836 60.02106683660.021\,066\,836\,60.021 066 836 6
HackerNews 0.007 869 082 705 826 4280.0078690827058264280.007\,869\,082\,705\,826\,4280.007 869 082 705 826 428 0.008 439 486 846 327 7820.0084394868463277820.008\,439\,486\,846\,327\,7820.008 439 486 846 327 782 0.034 344 911 20.03434491120.034\,344\,911\,20.034 344 911 2 0.047 427 612 870.047427612870.047\,427\,612\,870.047 427 612 87 0.032 764 945 207 651 0740.0327649452076510740.032\,764\,945\,207\,651\,0740.032 764 945 207 651 074 0.040 357 348 60.04035734860.040\,357\,348\,60.040 357 348 6
NIH ExPorter 0.004 653 726 876 369 9890.0046537268763699890.004\,653\,726\,876\,369\,9890.004 653 726 876 369 989 0.008 403 019 979 596 1380.0084030199795961380.008\,403\,019\,979\,596\,1380.008 403 019 979 596 138 0.043 679 299 90.04367929990.043\,679\,299\,90.043 679 299 9 0.040 626 581 40.04062658140.040\,626\,581\,40.040 626 581 4 0.035 694 541 166 612 130.035694541166612130.035\,694\,541\,166\,612\,130.035 694 541 166 612 13 0.032 072 096 30.03207209630.032\,072\,096\,30.032 072 096 3
OpenSubtitles 0.010 984 762 173 998 1670.0109847621739981670.010\,984\,762\,173\,998\,1670.010 984 762 173 998 167 0.003 219 955 600 798 130.003219955600798130.003\,219\,955\,600\,798\,130.003 219 955 600 798 13 0.016 057 877 90.01605787790.016\,057\,877\,90.016 057 877 9 0.020 182 878 80.02018287880.020\,182\,878\,80.020 182 878 8 0.006 519 081 320 969 960.006519081320969960.006\,519\,081\,320\,969\,960.006 519 081 320 969 96 0.020 155 954 80.02015595480.020\,155\,954\,80.020 155 954 8
OpenWebText2 0.124 183 724 740 787 390.124183724740787390.124\,183\,724\,740\,787\,390.124 183 724 740 787 39 0.190 496 414 899 826 050.190496414899826050.190\,496\,414\,899\,826\,050.190 496 414 899 826 05 0.071 949 544 40.07194954440.071\,949\,544\,40.071 949 544 4 0.076 104 828 90.07610482890.076\,104\,828\,90.076 104 828 9 0.110 142 648 978 081 850.110142648978081850.110\,142\,648\,978\,081\,850.110 142 648 978 081 85 0.051 656 098 10.05165609810.051\,656\,098\,10.051 656 098 1
PhilPapers 0.003 153 536 367 082 020 30.00315353636708202030.003\,153\,536\,367\,082\,020\,30.003 153 536 367 082 020 3 0.009 265 206 754 207 6110.0092652067542076110.009\,265\,206\,754\,207\,6110.009 265 206 754 207 611 0.077 201 461 80.07720146180.077\,201\,461\,80.077 201 461 8 0.060 257 032 10.06025703210.060\,257\,032\,10.060 257 032 1 0.093 572 527 923 277 930.093572527923277930.093\,572\,527\,923\,277\,930.093 572 527 923 277 93 0.039 030 510 90.03903051090.039\,030\,510\,90.039 030 510 9
Pile-CC 0.109 025 194 589 425 430.109025194589425430.109\,025\,194\,589\,425\,430.109 025 194 589 425 43 0.137 887 090 444 564 820.137887090444564820.137\,887\,090\,444\,564\,820.137 887 090 444 564 82 0.054 862 970 20.05486297020.054\,862\,970\,20.054 862 970 2 0.084 118 549 80.08411854980.084\,118\,549\,80.084 118 549 8 0.092 829 581 648 029 330.092829581648029330.092\,829\,581\,648\,029\,330.092 829 581 648 029 33 0.040 536 468 10.04053646810.040\,536\,468\,10.040 536 468 1
PubMed Abstracts 0.075 582 894 069 492 30.07558289406949230.075\,582\,894\,069\,492\,30.075 582 894 069 492 3 0.096 987 292 170 524 60.09698729217052460.096\,987\,292\,170\,524\,60.096 987 292 170 524 6 0.054 140 075 20.05414007520.054\,140\,075\,20.054 140 075 2 0.049 705 101 90.04970510190.049\,705\,101\,90.049 705 101 9 0.054 129 699 600 261 680.054129699600261680.054\,129\,699\,600\,261\,680.054 129 699 600 261 68 0.042 292 640 60.04229264060.042\,292\,640\,60.042 292 640 6
PubMed Central 0.113 906 043 489 830 370.113906043489830370.113\,906\,043\,489\,830\,370.113 906 043 489 830 37 0.060 811 433 941 125 870.060811433941125870.060\,811\,433\,941\,125\,870.060 811 433 941 125 87 0.053 594 405 80.05359440580.053\,594\,405\,80.053 594 405 8 0.046 177 139 20.04617713920.046\,177\,139\,20.046 177 139 2 0.049 780 788 423 504 3350.0497807884235043350.049\,780\,788\,423\,504\,3350.049 780 788 423 504 335 0.038 300 844 10.03830084410.038\,300\,844\,10.038 300 844 1
StackExchange 0.090 734 217 502 812 910.090734217502812910.090\,734\,217\,502\,812\,910.090 734 217 502 812 91 0.074 640 922 248 363 50.07464092224836350.074\,640\,922\,248\,363\,50.074 640 922 248 363 5 0.043 131 080 10.04313108010.043\,131\,080\,10.043 131 080 1 0.050 556 010 20.05055601020.050\,556\,010\,20.050 556 010 2 0.043 860 957 874 563 030.043860957874563030.043\,860\,957\,874\,563\,030.043 860 957 874 563 03 0.042 808 443 40.04280844340.042\,808\,443\,40.042 808 443 4
USPTO Backgrounds 0.040 149 048 461 867 9460.0401490484618679460.040\,149\,048\,461\,867\,9460.040 149 048 461 867 946 0.032 686 419 785 022 7360.0326864197850227360.032\,686\,419\,785\,022\,7360.032 686 419 785 022 736 0.036 183 449 80.03618344980.036\,183\,449\,80.036 183 449 8 0.042 328 963 60.04232896360.042\,328\,963\,60.042 328 963 6 0.030 807 951 127 831 2530.0308079511278312530.030\,807\,951\,127\,831\,2530.030 807 951 127 831 253 0.032 103 679 40.03210367940.032\,103\,679\,40.032 103 679 4
Ubuntu IRC 0.009 786 490 667 033 4840.0097864906670334840.009\,786\,490\,667\,033\,4840.009 786 490 667 033 484 0.008 286 269 381 642 3420.0082862693816423420.008\,286\,269\,381\,642\,3420.008 286 269 381 642 342 0.037 403 657 20.03740365720.037\,403\,657\,20.037 403 657 2 0.028 142 111 90.02814211190.028\,142\,111\,90.028 142 111 9 0.021 173 214 785 918 8050.0211732147859188050.021\,173\,214\,785\,918\,8050.021 173 214 785 918 805 0.052 948 081 40.05294808140.052\,948\,081\,40.052 948 081 4
Wikipedia (en) 0.089 434 136 415 032 20.08943413641503220.089\,434\,136\,415\,032\,20.089 434 136 415 032 2 0.106 829 889 118 671 420.106829889118671420.106\,829\,889\,118\,671\,420.106 829 889 118 671 42 0.062 576 9370.0625769370.062\,576\,9370.062 576 937 0.063 892 6250.0638926250.063\,892\,6250.063 892 625 0.080 422 933 364 482 70.08042293336448270.080\,422\,933\,364\,482\,70.080 422 933 364 482 7 0.078 383 412 70.07838341270.078\,383\,412\,70.078 383 412 7
YoutubeSubtitles 0.007 196 568 461 740 4710.0071965684617404710.007\,196\,568\,461\,740\,4710.007 196 568 461 740 471 0.011 665 397 323 668 0030.0116653973236680030.011\,665\,397\,323\,668\,0030.011 665 397 323 668 003 0.066 979 222 10.06697922210.066\,979\,222\,10.066 979 222 1 0.035 336 1830.0353361830.035\,336\,1830.035 336 183 0.047 607 492 695 126 0760.0476074926951260760.047\,607\,492\,695\,126\,0760.047 607 492 695 126 076 0.071 645 233 90.07164523390.071\,645\,233\,90.071 645 233 9
Table 5: Data mixtures on the SlimPajama dataset
Domain Baseline DoReMi SE CE JE VNE
ArXiv 0.045 807 080.045807080.045\,807\,080.045 807 08 0.021 336 648 613 214 4930.0213366486132144930.021\,336\,648\,613\,214\,4930.021 336 648 613 214 493 0.073 809 270.073809270.073\,809\,270.073 809 27 0.070 955 440.070955440.070\,955\,440.070 955 44 0.033 891 790.033891790.033\,891\,790.033 891 79 0.059 833 563 30.05983356330.059\,833\,563\,30.059 833 563 3
Books 0.042 026 350.042026350.042\,026\,350.042 026 35 0.042 031 083 256 006 240.042031083256006240.042\,031\,083\,256\,006\,240.042 031 083 256 006 24 0.147 381 860.147381860.147\,381\,860.147 381 86 0.166 936 910.166936910.166\,936\,910.166 936 91 0.159 218 780.159218780.159\,218\,780.159 218 78 0.069 957 295 10.06995729510.069\,957\,295\,10.069 957 295 1
C4 0.266 015 580.266015580.266\,015\,580.266 015 58 0.286 445 230 245 590 20.28644523024559020.286\,445\,230\,245\,590\,20.286 445 230 245 590 2 0.161 890 50.16189050.161\,890\,50.161 890 5 0.207 945 120.207945120.207\,945\,120.207 945 12 0.217 855 150.217855150.217\,855\,150.217 855 15 0.126 489 2570.1264892570.126\,489\,2570.126 489 257
CommonCrawl 0.520 302 490.520302490.520\,302\,490.520 302 49 0.551 770 925 521 850 60.55177092552185060.551\,770\,925\,521\,850\,60.551 770 925 521 850 6 0.186 319 130.186319130.186\,319\,130.186 319 13 0.217 504 270.217504270.217\,504\,270.217 504 27 0.262 254 530.262254530.262\,254\,530.262 254 53 0.093 263 536 60.09326353660.093\,263\,536\,60.093 263 536 6
GitHub 0.052 204 040.052204040.052\,204\,040.052 204 04 0.019 108 377 397 060 3940.0191083773970603940.019\,108\,377\,397\,060\,3940.019 108 377 397 060 394 0.095 199 80.09519980.095\,199\,80.095 199 8 0.090 233 290.090233290.090\,233\,290.090 233 29 0.055 590 50.05559050.055\,590\,50.055 590 5 0.231 397 179 10.23139717910.231\,397\,179\,10.231 397 179 1
StackExchange 0.033 704 920.033704920.033\,704\,920.033 704 92 0.029 153 479 263 186 4550.0291534792631864550.029\,153\,479\,263\,186\,4550.029 153 479 263 186 455 0.115 233 610.115233610.115\,233\,610.115 233 61 0.117 680 870.117680870.117\,680\,870.117 680 87 0.087 757 230.087757230.087\,757\,230.087 757 23 0.108 772 286 10.10877228610.108\,772\,286\,10.108 772 286 1
Wikpedia 0.039 939 540.039939540.039\,939\,540.039 939 54 0.050 133 459 270 000 460.050133459270000460.050\,133\,459\,270\,000\,460.050 133 459 270 000 46 0.220 165 840.220165840.220\,165\,840.220 165 84 0.128 744 10.12874410.128\,744\,10.128 744 1 0.183 432 020.183432020.183\,432\,020.183 432 02 0.310 286 882 80.31028688280.310\,286\,882\,80.310 286 882 8